While sitting on top of a 10m building you drop a chestnut,when it reaches a height of 2.5m you droo another chestnut.What initial speed should you give the second chestnut if they're both to reach at the same time?
1
Expert's answer
2020-11-05T10:07:45-0500

"h=v^2/(2g)\to v=\sqrt{2gh}=\sqrt{2\cdot9.81\cdot(10-2.5)}=12.13(m/s)"


"2.5=vt+gt^2/2\to 4.905t^2+12.13t-2.5=0\to"


"t=0.191(s)"


So,


"10=ut+gt^2/2\to u=\frac{10-gt^2/2}{t}=\frac{10-9.81 \cdot 0.191^2/2}{0.191}=51.42(m/s )". Answer